OpenStudy (anonymous):

find the volume 10cm 14.5cm12cm

OpenStudy (anonymous):

Volume = l*b*h = 10 * 14.5 * 12 = your answer...solve it in calculator

OpenStudy (anonymous):

is it 2262 or no

OpenStudy (anonymous):

are you finding the volume of a cuboid???

OpenStudy (anonymous):

i got its 1740

OpenStudy (anonymous):

finding the volume is different for different shapes..

OpenStudy (anonymous):

1740 right

OpenStudy (anonymous):

@kausarsalley i m sure he/she meant cuboid

OpenStudy (anonymous):

oh okay.....i was asking her to be sure.